OpenStudy (anonymous):

Joey is saving up for a down payment on a car. He plans to invest $2,000 at the end of every year for 3 years. If the interest rate on the account is 2.05% compounding annually, what is the present value of the investment? (I did this an got 1,881.82 but its not one of the answer choices, help!)

OpenStudy (anonymous):

Is one of the answers $6249.38?

OpenStudy (anonymous):

no ): the answers are $4,180.20 $5,762.15 $5,891.08 $5,978.76

OpenStudy (tkhunny):

Build it! I = 0.0205 v = 1/(1+i) = 0.9799118 \(2000(v + v^{2} + v^{3}) = 2000\dfrac{v-v^{4}}{1-v} = 2000(2.8810768832)\) If you learn how to build, you never will wonder again if you have the right formula.